OpenStudy (anonymous):

Is there an easy way to factor the following? x^3-27

OpenStudy (callisto):

Yes. Use the identity \(a^3-b^3 = (a-b)(a^2+ab+b^2)\)

OpenStudy (anonymous):

and if it's x^3+27?

OpenStudy (callisto):

\(a^3+b^3 = (a+b)(a^2-ab+b^2)\)

OpenStudy (anonymous):

Thanks!

OpenStudy (callisto):

Welcome :)